# LLM Multi-Agent Systems: Challenges and Open Problems

### v1

@@ Abstract
This paper explores existing works of multi-agent systems and identify challenges that remain inadequately addressed. By leveraging the diverse capabilities and roles of individual agents within a multi-agent system, these systems can tackle complex tasks through collaboration. We discuss optimizing task allocation, fostering robust reasoning through iterative debates, managing complex and layered context information, and enhancing memory management to support the intricate interactions within multi-agent systems. We also explore the potential application of multi-agent systems in blockchain systems to shed light on their future development and application in real-world distributed systems.
@@ 1 Introduction
This paper provides on a comprehensive exploration into multi-agent systems, offering a survey of the existing works while shedding light on the challenges and open problems in it. We study major components in multi-agent systems, including planning and memory storage, and address unique challenges posed by multi-agent systems, compared with single-agent systems. We also explore potential application of multi-agent systems in blockchain systems from two perspectives, including 1) utilizing multi-agent systems as tools, and 2) assigning an agent to each blockchain node to make it represent the user, such that the agent can can complete some tasks on behalf of the user in the blockchain network.
@@ 3.1 Global Planning
Designing effective work flow based on the agents’ specializations. Partitioning responsibilities and designing an effective work flow for agents is crucial for ensuring that the tasks for each agent are executable while meaningful and directly contributes to the overall objective in multi-agent systems. The biggest challenge lies in the following perspectives: 1) the partition of work flow should maximize the utilization of each agent’s unique capabilities, i.e., each agent can handle a part of the task that matches its capabilities and expertise; 2) each agent’s tasks must align with the overall goal; and 3) the design must understand and consider the context for the overall tasks as well as each agent. This requires a deep understanding of the task at hand and the specific strengths and limitations of each agent in the system.
@@
Game Theory. Game theory provides a well-structured framework for understanding strategic interactions in multi-agent systems, particularly for systems that involve complex interactions among agents such as debates or discussions. A crucial concept in game theory is equilibrium, e.g., Nash Equilibrium (Kreps, 1989) and Stackelberg Equilibrium (Von Stackelberg, 2010; Conitzer & Sandholm, 2006), that describes a state where, given the strategies of others, no agent benefits from unilaterally changing their strategy. Game theory has been applied in multi-agent systems, especially Stackelberg equilibrium (Gerstgrasser & Parkes, 2023; Harris et al., 2023), as the structure of Stackelberg equilibrium contains is a leader agent and multiple follower agents, and such hierarchical architectures are wildely considered in multi-agent systems. As an example, (Gerstgrasser & Parkes, 2023) designs a general multi-agent framework to identify Stackelberg Equilibrium in Markov games, and (Harris et al., 2023) extend the Stackelberg model to allow agents to consider external context information, such as traffic and weather, etc. However, some problems are still challenging in multi-agent systems, such as defining an appropriate payoff structure for both the collective strategy and individual agents based on the context of the overall tasks, and efficiently achieving equilibrium states. These unresolved issues highlight the ongoing need for refinement in the application of game theory to complex multi-agent scenarios.
@@ 3.2 Single-Agent Task Decomposition
1) Chain of Thoughts (CoT) (Wei et al., 2022) that transforms big tasks into step-by-step manageable tasks to represent interpretation of the agents’ reasoning (or thinking) process.
2) Multiple CoTs (Wang et al., 2022a) that explores multiple independent CoT reasoning paths and return the one with the best output.
3) Program-of-Thoughts (PoT) (Chen et al., 2022) that uses language models to generate text and programming language statements, and finally an answer.
4) Table-of-Thoughts (Tab-CoT) (Ziqi & Lu, 2023) that utilize a tabular-format for reasoning, enabling the complex reasoning process to be explicitly modelled in a highly structured manner.
5) Tree-of-Thoughts (ToT) (Yao et al., 2023; Long, 2023) that extends CoT by formulating a tree structure to explore multiple reasoning possibilities at each step. It enables generating new thoughts based on a given arbitrary thought and possibly backtracking from it.
6) Graph-of-Thoughts-Rationale (GoT-Rationale) (Besta et al., 2023) that explores an arbitrary graph to enable aggregating arbitrary thoughts into a new one and enhancing the thoughts using loops.
7) Rationale-Augmented Ensembles (Wang et al., 2022b) that automatically aggregate across diverse rationales to overcome the brittleness of performance to sub-optimal rationales.
@@
Consistency in Objectives. In multi-agent systems, consistency in objectives is maintained across various levels, i.e., from overall goals down to individual agent tasks and their decomposed tasks. Each agent must understand and effectively utilize the layered contexts while ensuring its task and the decomposed sub-tasks to remain aligned with the overall goals. (Harris et al., 2023) extends the Stackelberg model (Von Stackelberg, 2010; Conitzer & Sandholm, 2006) to enable agents to incorporate external context information, such as context (or insights) provided by other agents. However, aligning the complex context with the decomposed tasks during reasoning remains an unresolved issue.
@@ 4 Agent Memory and Information Retrieval
The memory in single-LLM agent system refers to the agent’s ability to record, manage, and utilize data, such as past historical queries and some external data sources, to help inference and enhance decision-making and reasoning (Yao et al., 2023; Park et al., 2023; Li & Qiu, 2023; Wang et al., 2023; Guo et al., 2023). While the memory in a single-LLM agent system primarily focuses on internal data management and utilization, a multi-agent system requires the agents to work collaboratively to complete some tasks, necessitating not only the individual memory capabilities of each agent but also a sophisticated mechanism for sharing, integrating, and managing information across the different agents, thus poses challenges to memory and information retrieval.
@@ 4.1 Classifications of Memory in Multi-agent Systems
  • External data storage (e.g., RAG): This is an emerging area in LLM research where models are integrated with external data storage like vector databases, such that the agents can access additional knowledge from these databases, enhancing their ability to ground and enrich their responses (Lewis et al., 2020). This allows the LLM to produce responses that are more informative, accurate, and highly relevant to the specific context of the query.
@@ 5.2 Blockchain Nodes as Agents
Smart Contract Management and Optimization. Smart contracts are programs that execute the terms of a contract between a buyer and a seller in a blockchain system. The codes are fixed, and are self-executed when predetermined conditions are met. Multi-agent systems can automate and optimize the execution of smart contracts with more flexible terms and even dynamic external information from users. Agents can negotiate contract terms on behalf of their users, manage contract execution, and even optimize gas fees (in the context of Ethereum (Wood et al., 2014). The agents can analyze context information , such as past actions and pre-defined criteria, and utilize the information with more flexibility. Such negotiations can also utilize game theory, such as Stackelberg Equilibrium (Von Stackelberg, 2010; Conitzer & Sandholm, 2006) when there is a leader negotiator and Nash Equilibrium (Kreps, 1989) when there is no leader.

### v2

@@ Abstract
This paper explores multi-agent systems and identify challenges that remain inadequately addressed. By leveraging the diverse capabilities and roles of individual agents, multi-agent systems can tackle complex tasks through agent collaboration. We discuss optimizing task allocation, fostering robust reasoning through iterative debates, managing complex and layered context information, and enhancing memory management to support the intricate interactions within multi-agent systems. We also explore potential applications of multi-agent systems in blockchain systems to shed light on their future development and application in real-world distributed systems.
@@ 1 Introduction
This paper explores multi-agent systems, offering a survey of the existing works while shedding light on the challenges and open problems in it. We study major components in multi-agent systems, including planning and memory storage, and address unique challenges posed by multi-agent systems, compared with single-agent systems. We also explore potential application of multi-agent systems in blockchain systems from two perspectives, including 1) utilizing multi-agent systems as tools, and 2) assigning an agent to each blockchain node to make it represent the user, such that the agent can can complete some tasks on behalf of the user in the blockchain network.
@@ 3.1 Global Planning
Designing effective work flow based on the agents’ specializations. Partitioning responsibilities and designing effective work flows for agents is crucial for ensuring that the tasks for each agent are executable while meaningful and directly contributes to the overall objective in systems. The biggest challenge lies in the following perspectives: 1) the partition of work flow should maximize the utilization of each agent’s unique capabilities, i.e., each agent can handle a part of the task that matches its capabilities and expertise; 2) each agent’s tasks must align with the overall goal; and 3) the design must understand and consider the context for the overall tasks as well as each agent. This requires a deep understanding of the task at hand and the specific strengths and limitations of each agent in the system.
@@
Game Theory. Game theory provides a well-structured framework for understanding strategic interactions in multi-agent systems, particularly for systems that involve complex interactions among agents such as debates or discussions. A crucial concept in game theory is equilibrium, e.g., Nash Equilibrium (Kreps, 1989) and Stackelberg Equilibrium (Von Stackelberg, 2010; Conitzer & Sandholm, 2006), that describes a state where, given the strategies of others, no agent benefits from unilaterally changing their strategy. Game theory has been applied in multi-agent systems, especially Stackelberg equilibrium (Gerstgrasser & Parkes, 2023; Harris et al., 2023), as the structure of Stackelberg equilibrium contains is a leader agent and multiple follower agents, and such hierarchical architectures are wildely considered in multi-agent systems. (Gerstgrasser & Parkes, 2023) designs a general multi-agent framework to identify Stackelberg Equilibrium in Markov games, and (Harris et al., 2023) extend the Stackelberg model to allow agents to consider external context information, such as traffic and weather, etc. However, some problems are still challenging in multi-agent systems, such as defining an appropriate payoff structure for both the collective strategy and individual agents based on the context of the overall tasks, and efficiently achieving equilibrium states. These unresolved issues highlight the ongoing need for refinement in the application of game theory to complex multi-agent scenarios.
@@ 3.2 Single-Agent Task Decomposition
i) Chain of Thoughts (CoT) (Wei et al., 2022) that transforms big tasks into step-by-step manageable tasks to represent interpretation of the agents’ reasoning (or thinking) process.
ii) Multiple CoTs (Wang et al., 2022a) that explores multiple independent CoT reasoning paths and return the one with the best output.
iii) Program-of-Thoughts (PoT) (Chen et al., 2022) that uses language models to generate text and programming language statements, and finally an answer.
iv) Table-of-Thoughts (Tab-CoT) (Ziqi & Lu, 2023) that utilize a tabular-format for reasoning, enabling the complex reasoning process to be explicitly modelled in a highly structured manner.
v) Tree-of-Thoughts (ToT) (Yao et al., 2023; Long, 2023) that extends CoT by formulating a tree structure to explore multiple reasoning possibilities at each step. It enables generating new thoughts based on a given arbitrary thought and possibly backtracking from it.
vi) Graph-of-Thoughts-Rationale (GoT-Rationale) (Besta et al., 2023) that explores an arbitrary graph to enable aggregating arbitrary thoughts into a new one and enhancing the thoughts using loops.
vii) Rationale-Augmented Ensembles (Wang et al., 2022b) that automatically aggregate across diverse rationales to overcome the brittleness of performance to sub-optimal rationales.
@@
Consistency in Objectives. In multi-agent systems, consistency in objectives is maintained across various levels, i.e., from overall goals down to individual agent tasks and their decomposed tasks. Each agent must understand and effectively utilize the layered contexts while ensuring its task and the decomposed sub-tasks to remain aligned with the overall goals. (Harris et al., 2023) extends the Stackelberg model (Von Stackelberg, 2010; Conitzer & Sandholm, 2006) to enable agents to incorporate external context information, such as context (or insights) provided by other agents. However, aligning the complex context with the decomposed tasks during reasoning remains unresolved.
@@ 4 Agent Memory and Information Retrieval
The memory in single-LLM agent systems refers to the agent’s ability to record, manage, and utilize data, such as past historical queries and some external data sources, to help inference and enhance decision-making and reasoning (Yao et al., 2023; Park et al., 2023; Li & Qiu, 2023; Wang et al., 2023; Guo et al., 2023). While the memory in a single-LLM agent system primarily focuses on internal data management and utilization, a multi-agent system requires agents to work collaboratively to complete some tasks, necessitating the individual memory capabilities of each agent as well as a sophisticated mechanism for sharing, integrating, and managing information across agents, thus poses challenges to memory and information retrieval.
@@ 4.1 Classifications of Memory in Multi-agent Systems
  • External data storage: This is an emerging area in LLM research where models are integrated with external data storage like vector databases, such that the agents can access additional knowledge from these databases, enhancing their ability to ground and enrich their responses (Lewis et al., 2020). This allows the LLM to produce responses that are more informative, accurate, and highly relevant to the specific context of the query.
@@ 5.2 Blockchain Nodes as Agents
Smart Contract Management and Optimization. Smart contracts are programs that execute the terms of a contract between a buyer and a seller in a blockchain system. The codes are fixed, and are self-executed when predetermined conditions are met. Multi-agent systems can automate and optimize the execution of smart contracts with more flexible terms and even dynamic external information from users. Agents can negotiate contract terms on behalf of their users, manage contract execution, and even optimize gas fees (in the context of Ethereum (Wood et al., 2014). The agents can analyze context information , such as past actions and pre-defined criteria, and utilize the information with flexibility. Such negotiations can also utilize game theory, such as Stackelberg Equilibrium (Von Stackelberg, 2010; Conitzer & Sandholm, 2006) when there is a leader negotiator and Nash Equilibrium (Kreps, 1989) when no leader exists.